OpenStudy (anonymous):

Why can we multiply \[\sqrt{x-1}*\sqrt{x+1}=sqrt{x²-1}\]? At first I thought it had to do with the domain but it is not the same... \[dom \sqrt{x-1}=[1;+\infty)\] \[dom \sqrt{x+1}=[-1;+\infty\] \[dom \sqrt{x²-1}=(-\infty;-1]U[1;+\infty)\]

OpenStudy (anonymous):

Fixed some displaying errors: Why can we multiply \[\sqrt{x-1}*\sqrt{x+1}=\sqrt{x²-1}\]? At first I thought it had to do with the domain but it is not the same... \[dom \sqrt{x-1}=[1;+\infty)\] \[dom \sqrt{x+1}=[-1;+\infty)\] \[dom \sqrt{x²-1}=(-\infty;-1]U[1;+\infty)\]

OpenStudy (anonymous):

ok so you can multiply but the domain will be \([1,\infty)\)

OpenStudy (anonymous):

Say I have example but the domain is mostly negative when multiplied (e.g. (-infty; 4]) does that mean I can still multiply but the domain for which the function is defined is [0;4]? I don't really get when I can multiply and when not.

OpenStudy (anonymous):

you can always multiply just have to restrict the domain

OpenStudy (anonymous):

Thank you